OpenStudy (anonymous):

The function f has first derivative given by f'(x)=(x)^1/2/1+x+x^3. What is the x-coordinate of the inflection point of the graph of f? a. 1.008 b. 0.473 c. 0 d. -0.278 e. the graph has no inflection points

OpenStudy (australopithecus):

Find the numbers not in the domain of, f'(x) Set the f'(x) = 0, and solve for x Create intervals out the numbers generated and input numbers inbetween these intervals. If you get a positive number out f'(x)>0, thus the function is increasing if you get a negative number out f'(x)<0, thus the function is decreasing an inflection point is likely when you have two intervals where the function is increasing or decreasing. They look like this 

OpenStudy (australopithecus):

Note that the intervals have to be next to eachother.

OpenStudy (anonymous):

could you show thiswork please?

OpenStudy (australopithecus):

for example a functin such as f'(x) = 3x/(x-1) we see the domain is All real numbers - (+1) so we have the number 1, we set f'(x) = 0 0(x-1) = 3x(x-1)/(x-1) 0 = 3x x = 0 we have a second number 0 so we have two numbers to generate intervals for 0, 1 we have the interval (-infinity, 0), (0,1), (1, infinity) we then make a table  Now insert numbers inbetween each interval into f'(x) so we will check (-infinity,0), lets use -1 as it belongs to this interval f'(-1) = 3(-1)/((-1-1) f'(-1) = -3/-2 = 3/2 Thus we have a positive the interval is increasing We know this because the rules: f'(x)>0, thus the function is increasing f'(x)<0, thus the function is decreasing Thus we have that